Find the value of:
a. log sin 62° 22' 33"
b. log cot 28° 13' 17"
c.log cos 125° 15' 23"
d.log tan 78° 45' 50"

Answer by Edwin McCravy(6938)


Find the value of:
a. log sin 62° 22' 33"

Make sure calculator is in degree mode.
Find 62 + 22/60 + 33/3600 =
Then take the sin of that
Then take the log of that

I get -.0525623137



b. log cot 28° 13' 17"

Make sure calculator is in degree mode.
Find 28 + 13/60 + 17/3600 =
Then take the tan of that
Then take the reciprocal of that.
Then take the log of that

I get .2702878049 


c.log cos 125° 15' 23"

There is no answer to this, but
let's try anyway and see why.

Make sure calculator is in degree mode.
Find 125 + 15/60 + 23/3600 =
Then take the cos of that
Then take the log of that

I get an ERROR message.  The reason is
because this is not defined. The angle is
in quadrant II, making the cosine negative.
However the log of a negative number is not
defined. So there is no answer to this.

d.log tan 78° 45' 50"
Make sure calculator is in degree mode.
Find 78 + 45/60 + 50/3600 =
Then take the tan of that
Then take the log of that
  
I get .7018887347   

Edwin